# Inequalities between Neumann and Dirichlet Laplacian eigenvalues on planar domains

### v2

@@ 1. Introduction
It has been conjectured that the inequality (1.3) should be true for arbitrary bounded domains in $\mathbb{R}^{d}$, see, e.g., [18, Conjecture 3.2.42]. However, the proofs by Payne and Levine and Weinberger make explicit use of the signs of curvatures of the boundary and do not extend to non-convex domains. A suspected counterexample on a non-convex domain proposed in [17, p. 207] based on numerical computations turned out not to hold; cf. [4, Remark 3.9]. On the other hand, a possible path towards proving (1.3) for more general domains was proposed in [4], but without actually proving it; see Remark 3.9 there. Very recently, the question and its significance have been pointed out again in [6].
@@
holds for all $k\in\mathbb{N}$, where $\omega(u)=\partial_{1}u_{2}-\partial_{2}u_{2}$ is the vorticity of a vector field $u$ and ${\mathcal{H}}_{\mathfrak{a}}$ consists of all vector fields $u\in L^{2}(\Omega)^{2}$ such that $\operatorname{div}u,\omega(u)\in L^{2}(\Omega)$ and $u|_{\partial\Omega}$ is tangential; see Theorem 3.4. After establishing this, we may essentially use vector fields $u$ whose components are eigenfunctions of $-\Delta_{\rm D}$ as test functions to obtain (1.4). A similar variational principle was derived recently by the author in connection with the hot spots conjecture [33]. These joint variational principles for Neumann and Dirichlet eigenvalues seem to be unique to the case $d=2$ and do not have any obvious analogs in higher space dimensions. Therefore it is likely that the approach taken here cannot be extended to higher dimensions.
@@ 2. Preliminaries
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded, connected Lipschitz domain; cf., e.g., [21, Chapter 3]. By Rademacher’s theorem, the unit normal vector $\nu(x)=(\nu_{1}(x),\nu_{2}(x))^{\top}$ and the unit tangent vector $\tau(x)=(\tau_{1}(x),\tau_{2}(x))^{\top}=(-\nu_{2}(x),\nu_{1}(x))^{\top}$ are uniquely defined for almost all $x\in\partial\Omega$. We denote by $H^{1}(\Omega)$ and $H^{2}(\Omega)$ the usual Sobolev spaces on $\Omega$ of order one and two, respectively. On the boundary $\partial\Omega$ we will make use of the Sobolev space $H^{1/2}(\partial\Omega)$ of order $1/2$ and its dual space $H^{-1/2}(\partial\Omega)$. We write $(\cdot,\cdot)_{\partial\Omega}$ for the sesquilinear duality between $H^{-1/2}(\partial\Omega)$ and $H^{1/2}(\partial\Omega)$. Recall that the trace map
@@
is well-defined by the above. For $u\in H^{2}(\Omega)$, $\partial_{\nu}\psi|_{\partial\Omega}$ can alternatively be defined by taking the trace $\nabla\psi|_{\partial\Omega}\in H^{1/2}(\partial\Omega)^{2}$.
@@
giving the statement of the lemma. ∎
@@ 3. A novel variational principle for Neumann and Dirichlet Laplacian eigenvalues
We only need to prove the mentioned properties of $\mathfrak{a}$. The characterization of $A$ and its domain follows from abstract theory, see, e.g., [29, Section X.3]. As $C_{0}^{\infty}(\Omega)^{2}\subset\mathrm{dom}\,\mathfrak{a}$, $\mathrm{dom}\,\mathfrak{a}$ is dense in $L^{2}(\Omega)^{2}$. Therefore it remains to show that $\mathfrak{a}$ is closed. To this end, let $(u^{n})$ be a Cauchy sequence in ${\mathcal{H}}_{\mathfrak{a}}$, that is,
@@
$\mathfrak{a}[w,v]=\int_{\Omega}\left(\operatorname{div}w\,\overline{\operatorname{div}v}+\omega(w)\overline{\omega(u)}\right)=0.$
@@
The statement of Proposition 3.2 on the spectrum of $A$ implies immediately the following variational principle; cf. [30, Section XIII.1]. Recall that
@@
where $A_{\rm N}$, $A_{\rm D}$ and $A_{0}$ are self-adjoint operators in $\nabla H^{1}(\Omega),\nabla^{\perp}H_{0}^{1}(\Omega)$ and $H_{\rm c}$, respectively. In particular, the spectrum of $A_{\rm N}$ equals the positive spectrum of $-\Delta_{\rm N}$, the spectrum of $A_{\rm D}$ equals the spectrum of $-\Delta_{\rm D}$, and $A_{\rm c}$ is the zero operator on $H_{\rm c}$.
@@
In fact, under those regularity assumptions, elliptic regularity implies that ${\mathcal{H}}_{\mathfrak{a}}={\mathcal{H}}_{\rm N}$, and a computation similar to the one in the proof of [33, Lemma 3.3] shows that the right-hand sides of (3.3) and (3.4) coincide.
@@
If $\partial\Omega$ is, e.g., piecewise $C^{1,1}$-regular and possible corners are convex, then ${\mathcal{H}}_{\mathfrak{a}}\subset H^{1}(\Omega)^{2}$; cf. the previous remark. In this case, ${\mathcal{H}}_{\mathfrak{a}}$ is compactly embedded into $L^{2}(\Omega)^{2}$ and, hence, $A$ has a purely discrete spectrum, implying that ${\mathcal{H}}_{\rm c}=\ker A$ is finite-dimensional. However, it can in fact be shown that $\dim H_{\rm c}$ equals the first Betti number of $\Omega$, see, e.g., [22, Chapter 3].
@@
for all $k\in\mathbb{N}$. However, this principle is not sufficient for the argument carried out in the proof of the main result of this article below, as the test functions used there are no gradient fields, i.e., they do not lie in the required test space.
@@ 4. Dirichlet-Neumann eigenvalue inequalities on simply connected domains
The inequality (4.3) follows from Theorem 4.1 and the fact that the lowest eigenvalue of $-\Delta_{\rm D}$ is always simple. Furthermore, closed nodal lines can be excluded by the following standard argument: let $\psi$ be an eigenfunction of $-\Delta_{\rm D}$ corresponding to $\mu\in\{\mu_{2},\mu_{3}\}$ and assume, for a contradiction, that $\psi$ has a closed nodal line ${\mathcal{C}}$; let $\Omega_{0}\subset\Omega$ be the open set for which $\partial\Omega_{0}={\mathcal{C}}$. Then $\psi|_{\Omega_{0}}$ is an eigenfunction of the Laplacian on $\Omega_{0}$ with Dirichlet boundary conditions. Using domain monotonicity of Dirichlet Laplacian eigenvalues, we obtain, denoting by $\lambda_{1}(\Omega_{0})$ the smallest eigenvalue of the Dirichlet Laplacian on $\Omega_{0}$,
@@
Before we proceed to the proof of Theorem 4.1, we provide the following lemma, see, e.g., [24, Lemma 2.2]. Its proof is based on a unique continuation argument.
@@
To prove the inequalities of the theorem, fix $k\in\mathbb{N}$ and let $\varphi_{j}$, $j=1,\dots,k$, be an orthonormal set of eigenfunctions of $-\Delta_{\rm D}$ such that $-\Delta_{\rm D}\varphi_{j}=\lambda\varphi_{j}$ holds for $j=1,\dots,k$. Define
@@
Indeed, the vector fields $u$ of the form (4.4) span a $2k$-dimensional subspace of $\mathrm{dom}\,\mathfrak{a}$, and the fields $v=\nabla^{\perp}\varphi$ with $\varphi\in\ker(-\Delta_{\rm D}-\lambda_{k})$ span a subspace of $\ker(A-\lambda_{k})$ of dimension $\dim\ker(-\Delta_{\rm D}-\lambda_{k})$. Moreover, assume there exists $u$ belonging to both spaces, i.e. $u$ can be written (4.4) and $u=\nabla^{\perp}\varphi$ for some $\varphi\in\ker(-\Delta_{\rm D}-\lambda_{k})$. Then $\varphi$ belongs to $H^{2}(\Omega)$, and
@@
Now we verify (4.8) under the conditions for strict inequality specified in the theorem. Consider first the case where $\lambda_{k}$ is a simple eigenvalue of $-\Delta_{\rm D}$. Assume that $u$ is of the form (4.4) and, at the same time belongs to $\ker(A-\lambda_{k})$. Based on the differential equation $Au=\lambda_{k}u$ and the structure of the eigenspaces of $A$ analyzed in Proposition 3.2 this can be written
@@
Thus, $\varphi_{k}\in H^{2}(\Omega)$, and decomposition of the trace of $\nabla\varphi_{k}$ into its normal and tangential components gives
@@
where we have employed (4.9) and Lemma 2.2. Note that there exists a relatively open, non-empty set $\omega\subset\partial\Omega$ such that $(\alpha,\beta)^{\top}$ is nowhere normal on $\omega$. Hence, (4.10) yields $\partial_{\nu}\varphi_{k}|_{\omega}=0$, and since $\varphi_{k}|_{\partial\Omega}=0$, Lemma 4.4 implies $\varphi_{k}=0$ constantly in $\Omega$, a contradiction. This gives (4.8).
@@
Note that there exists a non-empty, open set $\Sigma^{\prime}\subset\Sigma$ such that $\overline{\Sigma^{\prime}}\subset\Sigma$ on which $u$ is smooth up to the boundary; cf., e.g., [21, Theorem 4.18 (ii)]. Moreover,
@@
Let $\Omega$ be the unit disk in $\mathbb{R}^{2}$. Then the eigenvalues of $-\Delta_{\rm D}$ are the squared zeroes of the Bessel functions $J_{0}$ (with multiplicity one) and $J_{1},J_{2},\dots$ (with multiplicity two). On the other hand, the positive eigenvalues of $-\Delta_{\rm N}$ are the squares of the positive roots of the derivatives $J_{0}^{\prime}$ (with multiplicity one) and $J_{1}^{\prime},J_{2}^{\prime},\dots$ (with multiplicity two); see, e.g., [26, Section 9.5.3]. Concretely,

### v4

@@ 1. Introduction
It has been conjectured that the inequality (1.3) should be true for arbitrary bounded domains in $\mathbb{R}^{d}$, see, e.g., [20, Conjecture 3.2.42]. However, the proofs by Payne and Levine and Weinberger make explicit use of the signs of curvatures of the boundary and do not extend to non-convex domains. A suspected counterexample on a non-convex domain proposed in [19, p. 207] based on numerical computations turned out not to hold; cf. [4, Remark 3.9]. On the other hand, a possible path towards proving (1.3) for more general domains was proposed in [4], but without actually proving it; see Remark 3.9 there. Very recently, the question and its significance have been pointed out again in [6].
@@
holds for all $k\in\mathbb{N}$, where $\omega(u)=\partial_{1}u_{2}-\partial_{2}u_{1}$ is the vorticity of a vector field $u$ and ${\mathcal{H}}_{\mathfrak{a}}$ consists of all vector fields $u\in L^{2}(\Omega)^{2}$ such that $\operatorname{div}u,\omega(u)\in L^{2}(\Omega)$ and $u|_{\partial\Omega}$ is tangential; see Theorem 3.4. After establishing this, we may essentially use vector fields $u$ whose components are eigenfunctions of $-\Delta_{\rm D}$ as test functions to obtain (1.4). A similar variational principle was derived recently by the author in connection with the hot spots conjecture [36]. These joint variational principles for Neumann and Dirichlet eigenvalues seem to be unique to the case $d=2$ and do not have any obvious analogs in higher space dimensions. Therefore it is likely that the approach taken here cannot be extended to higher dimensions.
@@ 2. Preliminaries
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded, connected Lipschitz domain; cf., e.g., [23, Chapter 3]. By Rademacher’s theorem, the unit normal vector $\nu(x)=(\nu_{1}(x),\nu_{2}(x))^{\top}$ and the unit tangent vector $\tau(x)=(\tau_{1}(x),\tau_{2}(x))^{\top}=(-\nu_{2}(x),\nu_{1}(x))^{\top}$ are uniquely defined for almost all $x\in\partial\Omega$. We denote by $H^{1}(\Omega)$ and $H^{2}(\Omega)$ the usual Sobolev spaces on $\Omega$ of order one and two, respectively. On the boundary $\partial\Omega$ we will make use of the Sobolev space $H^{1/2}(\partial\Omega)$ of order $1/2$ and its dual space $H^{-1/2}(\partial\Omega)$. We write $(\cdot,\cdot)_{\partial\Omega}$ for the sesquilinear duality between $H^{-1/2}(\partial\Omega)$ and $H^{1/2}(\partial\Omega)$. Recall that the trace map
@@
is well-defined by the above. For $u\in H^{2}(\Omega)$, $\partial_{\nu}\psi|_{\partial\Omega}$ can alternatively be defined by taking the trace $\nabla\psi|_{\partial\Omega}\in H^{1/2}(\partial\Omega)^{2}$ (though even in this case $\partial_{\nu}\psi$ does not necessarily belong to $H^{1/2}(\partial\Omega)$ as $\nu$ in general is merely bounded).
@@
by the boundedness of the normal trace on ${\mathcal{E}}(\Omega)$, giving the statement of the lemma. ∎
@@ 3. A novel variational principle for Neumann and Dirichlet Laplacian eigenvalues
We only need to prove the mentioned properties of $\mathfrak{a}$. The characterization of $A$ and its domain follows from abstract theory, see, e.g., [32, Section X.3]. As $C_{0}^{\infty}(\Omega)^{2}\subset\mathrm{dom}\,\mathfrak{a}$, $\mathrm{dom}\,\mathfrak{a}$ is dense in $L^{2}(\Omega)^{2}$. Therefore it remains to show that $\mathfrak{a}$ is closed. To this end, let $(u^{n})$ be a Cauchy sequence in ${\mathcal{H}}_{\mathfrak{a}}$, that is,
@@
$\mathfrak{a}[w,v]=\int_{\Omega}\left(\operatorname{div}w\,\overline{\operatorname{div}v}+\omega(w)\overline{\omega(v)}\right)=0.$
@@
The statement of Proposition 3.2 on the spectrum of $A$ implies immediately the following variational principle; cf. [33, Section XIII.1]. Recall that
@@
where $A_{\rm N}$, $A_{\rm D}$ and $A_{\rm c}$ are self-adjoint operators in $\nabla H^{1}(\Omega),\nabla^{\perp}H_{0}^{1}(\Omega)$ and $H_{\rm c}$, respectively. In particular, the spectrum of $A_{\rm N}$ equals the positive spectrum of $-\Delta_{\rm N}$, the spectrum of $A_{\rm D}$ equals the spectrum of $-\Delta_{\rm D}$, and $A_{\rm c}$ is the zero operator on $H_{\rm c}$.
@@
In fact, under those regularity assumptions, elliptic regularity implies that ${\mathcal{H}}_{\mathfrak{a}}={\mathcal{H}}_{\rm N}$, and a computation similar to the one in the proof of [36, Lemma 3.3] shows that the right-hand sides of (3.3) and (3.4) coincide.
@@
If $\partial\Omega$ is, e.g., piecewise $C^{1,1}$-regular and possible corners are convex, then ${\mathcal{H}}_{\mathfrak{a}}\subset H^{1}(\Omega)^{2}$; cf. the previous remark. In this case, ${\mathcal{H}}_{\mathfrak{a}}$ is compactly embedded into $L^{2}(\Omega)^{2}$ and, hence, $A$ has a purely discrete spectrum, implying that ${\mathcal{H}}_{\rm c}=\ker A$ is finite-dimensional. However, it can in fact be shown that $\dim H_{\rm c}$ equals the first Betti number of $\Omega$, see, e.g., [24, Chapter 3].
@@
for all $k\in\mathbb{N}$. However, this principle is not sufficient for the argument carried out in the proof of the main result of this article below, as the test functions used there are not gradient fields, i.e., they do not lie in the required test space.
Remark 3.9.
The operator $A$ constructed in this section has a natural interpretation in the language of differential forms, which the author was not aware of initially. After a preprint version of the present manuscript had come out, this interpretation was worked out independently by Fries, Goffeng and Miranda [12], Hua, Münch and Zhang [15], and Muravyev [25]; the manuscript [15] also provides a generalization to a class of surfaces with boundary. In fact, upon identifying classical vector fields with 1-forms, the identity (2.3) corresponds to the Hodge decomposition, and the operator $A$ is the Hodge Laplacian with absolute boundary conditions. We refer to the mentioned manuscripts for more details.
@@ 4. Dirichlet-Neumann eigenvalue inequalities on simply connected domains
The inequality (4.3) follows from Theorem 4.1 and the fact that the lowest eigenvalue of $-\Delta_{\rm D}$ is always simple. Furthermore, closed nodal lines can be excluded by the following standard argument: let $\psi$ be an eigenfunction of $-\Delta_{\rm N}$ corresponding to $\mu\in\{\mu_{2},\mu_{3}\}$ and assume, for a contradiction, that $\psi$ has a closed nodal line ${\mathcal{C}}$; let $\Omega_{0}\subset\Omega$ be the open set for which $\partial\Omega_{0}={\mathcal{C}}$. Then $\psi|_{\Omega_{0}}$ is an eigenfunction of the Laplacian on $\Omega_{0}$ with Dirichlet boundary conditions. Using domain monotonicity of Dirichlet Laplacian eigenvalues, we obtain, denoting by $\lambda_{1}(\Omega_{0})$ the smallest eigenvalue of the Dirichlet Laplacian on $\Omega_{0}$,
@@
Before we proceed to the proof of Theorem 4.1, we provide the following lemma, see, e.g., [27, Lemma 2.2]. Its proof is based on a unique continuation argument.
@@
To prove the inequalities of the theorem, fix $k\in\mathbb{N}$ and let $\varphi_{j}$, $j=1,\dots,k$, be an orthonormal set of eigenfunctions of $-\Delta_{\rm D}$ such that $-\Delta_{\rm D}\varphi_{j}=\lambda_{j}\varphi_{j}$ holds for $j=1,\dots,k$. Define
@@
Indeed, the vector fields $u$ of the form (4.4) span a $2k$-dimensional subspace of $\mathrm{dom}\,\mathfrak{a}$, and the fields $v=\nabla^{\perp}\varphi$ with $\varphi\in\ker(-\Delta_{\rm D}-\lambda_{k})$ span a subspace of $\ker(A-\lambda_{k})$ of dimension $\dim\ker(-\Delta_{\rm D}-\lambda_{k})$. Moreover, assume there exists $u$ belonging to both spaces, i.e. $u$ can be written (4.4) and $u=\nabla^{\perp}\varphi$ for some $\varphi\in\ker(-\Delta_{\rm D}-\lambda_{k})$. Then $\varphi$ belongs to $H^{2}(\Omega)$ (since $\nabla^{\perp}\varphi=u\in H^{1}(\Omega)^{2}$), and
@@
Now we verify (4.8) under the conditions for strict inequality specified in the theorem. Consider first the case where $\lambda_{k}$ is a simple eigenvalue of $-\Delta_{\rm D}$. Assume that $u$ is of the form (4.4) and, at the same time, belongs to $\ker(A-\lambda_{k})$. Based on the differential equation $Au=\lambda_{k}u$ and the structure of the eigenspaces of $A$ analyzed in Proposition 3.2 this can be written
@@
Thus, $\nabla\varphi_{k}\in H^{1}(\Omega)^{2}$ and, hence, $\varphi_{k}\in H^{2}(\Omega)$. Moreover, decomposition of the trace of $\nabla\varphi_{k}$ into its normal and tangential components gives
@@
where we have employed (4.9) and Lemma 2.2. Note that there exists a relatively open, non-empty set $\Gamma\subset\partial\Omega$ such that $(\overline{\alpha},\overline{\beta})^{\top}$ and $\nu(x)$ are linearly independent for almost all $x\in\Gamma$. Hence, (4.10) yields $\partial_{\nu}\varphi_{k}|_{\Gamma}=0$, and since $\varphi_{k}|_{\partial\Omega}=0$, Lemma 4.4 implies $\varphi_{k}=0$ constantly in $\Omega$, a contradiction. This gives (4.8).
@@
Note that there exists a non-empty, open set $\Sigma^{\prime}\subset\Sigma$ such that $\overline{\Sigma^{\prime}}\subset\Sigma$ on which $u$ is smooth up to the boundary; cf., e.g., [23, Theorem 4.18 (ii)]. Moreover,
@@
Let $\Omega$ be the unit disk in $\mathbb{R}^{2}$. Then the eigenvalues of $-\Delta_{\rm D}$ are the squared zeroes of the Bessel functions $J_{0}$ (with multiplicity one) and $J_{1},J_{2},\dots$ (with multiplicity two). On the other hand, the positive eigenvalues of $-\Delta_{\rm N}$ are the squares of the positive roots of the derivatives $J_{0}^{\prime}$ (with multiplicity one) and $J_{1}^{\prime},J_{2}^{\prime},\dots$ (with multiplicity two); see, e.g., [29, Section 9.5.3]. Concretely,
@@ Acknowledgements
Data availability statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest statement
There is no conflict of interest.